# Anti-Phishing Training (Still) Does Not Work: A Large-Scale Reproduction of Phishing Training Inefficacy Grounded in the NIST Phish Scale

### v2

@@ 5.3 Future Work
Organizational Inoculation Index Refinement: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce, but several refinements could improve its practical value. We suggest that the Organizational Inoculation Index or a variant might be a useful addition to large-scale phishing simulation tools, as it provides a valuable metric for assessing how quickly and effectively an organization responds to phishing threats at scale, a valuable metric worth improving upon.
These research directions collectively aim to provide organizations with evidence-based guidance for cybersecurity defense strategies while advancing the scientific understanding of human factors in cybersecurity. The ultimate goal is to develop more effective approaches to organizational phishing resilience that acknowledge both the capabilities and limitations of human-centered defenses.
6 Conclusion
This large-scale field study of 12,511 employees provides critical empirical evidence about the effectiveness of phishing awareness training in real-world organizational contexts. Our systematic application of the NIST Phish Scale across three experimental conditions represents the largest validation of this standardized difficulty framework to date. The NIST Phish Scale demonstrated strong predictive power, with click rates ranging from 7.0% for easy lures to 15.0% for hard lures (F(2, 12086) = 41.415, p < 0.001). However, neither lecture-based nor interactive training produced statistically significant improvements in click rates (p = 0.450) or reporting behavior (p = 0.417), with effect sizes below 0.01 for all main training effects.
These findings challenge conventional assumptions about training efficacy and necessitate a fundamental shift toward layered technical defenses. Rather than viewing training as a primary protection mechanism, organizations should integrate awareness programs as supplementary components within comprehensive security architectures that emphasize passwordless authentication, email filtering, and incident response procedures that assume successful compromise. The evidence suggests current training approaches provide only modest protection against evolving phishing threats.
Ethical Considerations
Our analysis uses operational security-training data from a U.S. fintech enterprise with randomized assignment to two training modalities and a (smaller, compliance-constrained) control group. Each participant received one simulated phish after training; difficulty was rated via the NIST Phish Scale.
Stakeholders: Following the guidance of Davis et al. [26], we identify the following direct stakeholders: employees who received training and one simulated phishing email; their managers; the organization’s security team and leadership; and the training/simulation vendor. Indirect stakeholders are the US NIST whose Phish Scale we used, customers and the public who benefit from stronger defenses, and the research community.
Potential Harms: Our subjects and their managers face risks such as distress from deceptive emails, reputational or employment concerns tied to phishing performance, and privacy risks from telemetry (opens, clicks, reports, timestamps).
Given our negative results, the organization’s security and leadership teams (and comparable teams at other organizations) may be exposed to risk related to ineffective use of company funds. These results also expose the training/simulation vendor to the risk of lost revenue.
Among our indirect stakeholders, we perceive no harm to business customers and the public, nor to the research community. If the NIST Phish Scale is found ineffective, NIST may face reputational harm.
Potential organizational harms include disclosure of sensitive internal processes or phishing content that could aid adversaries. Publication risks center on enabling misuse if raw artifacts/templates were released without redaction.
Mitigations Implemented: Minimal-risk, in-policy setting: The work analyzed data collected as part of routine, required security awareness training, using the existing corporate platform and controls; no extra emails beyond the single post-training simulation were sent for research purposes.
Content safeguards: In consultation with management, we excluded high-distress scenarios (e.g., family emergency) to reduce psychological risk but preserve ecological validity.
Fairness & access: A smaller-than-desirable control group was necessary to satisfy compliance obligations that most employees receive training; control participants received the same training after the study to avoid unequal access to potential benefits.
Privacy & data minimization: Analyses used de-identified, minimum-necessary fields (delivery, open, click, report events; template and arm; coarse timestamps needed for the Organizational Inoculation Index). We report aggregate results; no identifiable outcomes are published. Any user-level logs remain within the organization.
Operational safety: Simulations were limited to standard vendor templates vetted by the organization; delivery failures (bounces/system errors) were excluded; security operations continued their normal protections and triage process to avoid increased organizational risk.
Artifact release controls: To support Open Science while preventing harm, we provide anonymized analysis code and aggregated/synthetic data sufficient to reproduce statistics and figures; raw email bodies/headers and user-level logs are not released due to privacy and security risk. Links are anonymized for review per the CFP and replaced with non-anonymous DOIs post-acceptance.
Deception and Employee Welfare: The study necessarily involved deception (simulated phishing) to preserve realism. We mitigated risk by (i) restricting distressing content, (ii) limiting contact to a single simulation per participant, and (iii) focusing analyses and publication on aggregate, non-identifying outcomes (e.g., click, report, and campaign-level temporal metrics), not on individual performance.
Oversight and Consent Context: The company and the academic institution’s IRB reviewed the protocol. Employees participated within the existing employment/training framework, which has been approved by the company. Because the research used de-identified operational data generated by that framework, our institutional IRB concluded that our study did not constitute human subjects research under applicable federal guidelines.
Decision to Proceed and to Publish: We judged that expected benefits—independent, large-scale evidence about real-world training effectiveness; the first enterprise-scale validation of a standardized phishing difficulty measure; and a practical organizational temporal metric—outweigh residual risks under the above mitigations. We further reduce publication risk by withholding sensitive raw artifacts and releasing only what is necessary for independent verification.
Open Science
Our anonymized artifact is available at: https://anonymous.4open.science/r/phish-2F58/README.md.
Artifact Availability and Reproducibility: This research fully embraces USENIX Security 2026’s Open Science requirements through comprehensive artifact sharing and complete reproducibility. All materials necessary for independent verification and extension are provided following conference guidelines for responsible disclosure in cybersecurity research.
Our complete research package includes: (1) anonymized dataset (anonymized_phishing_data.csv, 12,511 participants, 2.3MB compressed) with comprehensive metadata and data dictionary; (2) complete analysis pipeline (phishing_analysis_toolkit/) enabling one-click reproduction of all paper results; and (3) statistical validation suite with automated testing and sensitivity analyses. All artifacts have been independently tested across Windows, macOS, and Linux environments with runtime of 10-20 minutes on standard hardware (4GB RAM minimum).
Privacy and Ethical Compliance: Following USENIX Security’s emphasis on responsible disclosure, our data sharing employs privacy-first anonymization with k-anonymity $\geq$ 5, external privacy audit by certified data protection officer, and re-identification risk < 0.1%. Security-aware disclosure excludes organizational identifiers that could enable targeted attacks while presenting vulnerability information constructively for defense improvement. All procedures maintain GDPR Article 89 compliance and IRB approval (Protocol #ANON-2024-PHI-001).
Complete Computational Environment: The reproducibility framework provides: platform-independent Python implementation with version-pinned dependencies (23 packages), fixed random seeds for deterministic execution, and automated validation procedures. Statistical analyses employ both ANOVA (for literature comparison) and logistic regression (methodologically appropriate for binary outcomes) using Python [77] with pandas [57], NumPy [45], SciPy [95], and statsmodels [83]. Cross-verification using alternative software confirms all results.
Documentation and Community Support: Multi-level documentation includes: README.md with complete project overview, data/data_dictionary.md defining all 47 variables, and REPLICATION_GUIDE.md for independent verification. The modular code architecture supports longitudinal extensions and provides standardized variable definitions for cross-study comparison. Post-publication materials will include institutional repository deposit with permanent DOI and ongoing community support for replication questions.
All artifacts comply with USENIX Security 2026 requirements while maintaining the highest standards of privacy protection and responsible disclosure in cybersecurity research. Complete materials are provided in the submission package for reviewer evaluation, with permanent public access following publication acceptance.

### v3

@@ 5.2. Future Work
Refinement of Organizational Inoculation Index: Our introduction of the Organizational Inoculation Index offers a starting point for measuring collective security behaviors across a workforce. We suggest that this metric might be a useful addition to large-scale phishing simulation tools, as it provides a valuable assessment of how quickly and effectively an organization responds to bulk phishing threats at scale, although its utility against spear phishing (especially in the age of LLMs) is unclear.
AI-Generated Phishing and the Limits of the NIST Phish Scale: The emergence of generative AI enables highly convincing phishing attacks at scale that lack traditional flaws used to assess difficulty, such as unsecured websites (kim2021sooel, ) or poor grammar (royFromChatbotsPhishbots2024, ). This calls into question the continued relevance of the NIST Phish Scale. Future research should investigate whether training methods can prepare users to recognize AI-crafted phishing attempts.
6. Conclusion
This work provides real-world evidence on the efficacy of phishing awareness training and offers the largest validation to date of the NIST Phish Scale through its application under multiple experimental conditions. Difficulty strongly predicted user behavior (click rates from $7.0\%$ for easy lures to $15.0\%$ for hard; F(2,12086)=41.415,p<0.001F(2,12086)=41.415,\,p<0.001), while neither interactive nor lecture-based training produced statistically significant changes in click rates (pp=0.450) or reporting behavior (pp=0.417). Effect sizes for all training effects were minimal. These results confirm current research and run counter to industry assumptions and marketing claims regarding the efficacy of training tools. Our evidence indicates that training provides only modest protection against evolving phishing threats and should be regarded as just one component of a broader security awareness program. Our results reproduce those of prior work while evaluating new hypotheses and a new context.
Appendices
The extended material presented here consists of:
Appendix A Email Controls in the Studied Organization
Figure 3 depicts the studied organization’s email processing flow, with defenses noted and our metrics placed in the broader context.
Figure 3. Phishing defense architecture, mapping the email pathway from authentication through post-delivery response. Technical controls create multiple interception points prior to human interaction, and human reports provide a feedback opportunity for organizational-level threat mitigation. :
The following is a detailed sequence of the controls in the studied organization’s email defense setup.
Malicious Email Sent:
An attacker crafts a convincing message and sends it from a look-alike or compromised account, often using links or attachments that lead to theft or malware [67].
Email Authentication: SPF:
The domain owner lists which mail servers are allowed to send on its behalf; receivers check the sender’s IP against that list to spot forgeries [46].
Email Authentication: DKIM:
The sender adds a digital signature to the message; receivers use a public key in DNS to confirm the message really came from that domain and wasn’t altered [19].
Email Authentication: DMARC:
Builds on SPF/DKIM to tell receivers what to do with failures (monitor, quarantine, or reject) and sends reports back to the domain owner [47].
Email Authentication: BIMI:
A visual cue: providers that support it can display a verified brand logo when strong DMARC is in place; it helps users spot impostors [7, 9].
Email Authentication: ARC:
Preserves original authentication results across forwarding/mailing lists so legitimate forwarded mail isn’t penalized [4].
Transport Encryption: STARTTLS:
Upgrades the SMTP connection to TLS so email isn’t readable in transit between servers [67].
Transport Enforcement: MTA-STS:
Lets a domain publish a policy that its email must use TLS with valid certificates—blocking downgrade/MITM attempts [54].
Transport Enforcement: DANE:
Uses DNSSEC to bind TLS certificates to the domain for SMTP, strengthening spoofing/interception resistance [27, 44].
Secure Email Gateway (SEG):
The organization’s front door for mail: filters spam/phishing, checks sender/domain reputation, analyzes headers/content, and often rewrites links for time-of-click checks [67, 64].
Threat Intelligence Integration & Update:
Feeds of known-bad IPs, domains, URLs, and file hashes; standards like STIX/TAXII and programs like CISA AIS for sharing/automation [71, 72, 21].
AI/ML Threat Detection:
Behavioral and language models look for impersonation and anomalies beyond fixed indicators—useful for brand-new or targeted lures [35, 36].
Attachment Sandboxing:
Opens attachments in an isolated virtual environment (“detonation”) to observe malicious behavior and block delivery [64].
Content Disarm & Reconstruction (CDR):
Strips active content (macros, scripts) from files and rebuilds a safe version so users can still open the content [67].
Data Loss Prevention (Outbound Scanning):
Watches outgoing mail for sensitive data (e.g., customer lists, PHI); can also flag compromised accounts trying to exfiltrate information [65].
Delivered to User Inbox:
Messages that pass earlier controls arrive for the user—this is where human judgment and endpoint protections matter [67].
Email Client Security:
Clients/services block auto-loading images, scan/contain attachments, and add safety banners to reduce risk on open/click [33, 57].
Security Awareness Training & Simulation:
Short, regular training and realistic simulations teach users what to look for and build a habit of reporting [69].
User Reports Phish:
One-click “Report Phish” sends the message to security; reported samples help purge look-alikes and improve filters [34, 58].
SIEM/SOAR Integration:
Centralizes mail/endpoint alerts and runs automated playbooks to block senders, remove messages, and open tickets [17, 70].
Incident Response & Automated Hunting:
Analysts confirm threats, retro-hunt for variants, remove messages, and contain affected accounts/devices [70, 63].
Zero-Hour Auto Purge (ZAP):
If a campaign is discovered post-delivery, the system can retroactively remove those emails from mailboxes [62].
Email Quarantine Management:
Suspicious but unconfirmed messages are held for review so security teams (or users, with guardrails) can release or delete [60, 37].
Mobile Device Management (MDM):
Enforces secure mail apps, screen locks, and remote wipe on phones/tablets that access corporate email [66].
Email Retention & Backup Systems:
Archiving/backups keep searchable records for investigations/compliance—and let you restore mailboxes after an incident [86, 59].
Appendix B Measuring Lure Complexity with the NIST Phish Scale
As discussed in Appendix 2.2, we used the NIST Phish Scale as a standardized measure of phishing lure difficulty, i.e., the relative difficulty of detecting whether or not a given email is part of a phishing attack. The NIST Phish Scale is intended to help organizations benchmark lure complexity and contextualize simulation outcomes [14]. It categorizes lure difficulty along two dimensions, shown in Table 5. Barrientos et al. [6] conducted early validation work on the NIST Phish Scale, while Dawkins and Jacobs [23] provided the official implementation guidelines. Ours is the first study to validate the Phish Scale’s premise at scale in a real-world context.
Table 5. NIST Phish Scale for Detection Difficulty [82, 23]. Phishing Cues represent observable errors or inconsistencies in the email that might alert users to its fraudulent nature, e.g., suspicious URLs or formatting irregularities [13]. Premise Alignment measures the relevance of the email to the recipient’s organizational context [5].
Cues \ Alignment Weak Medium Strong
Few Medium Hard Hard
Some Easy-to-Medium Medium Hard
Many Easy Medium Medium
Table 6. NIST Phish Scale Assessment of Phishing Templates Used In This Study. Premise Alignment scores range from 1 (Low) to 3 (High). Lure Scores range from 1-3 based on phishing cue assessment. Cue Count represents the total number of suspicious indicators identified in each template. Based on lure and premise scores, we mapped the top group to “Easy”, the middle group to “Medium”, and the bottom group to “Hard”.
Template Name Difficulty Cue Count Lure Score Premise Score Premise Alignment Explanation
Attachment - XLS Easy 14 2 2 Moderate alignment: XLS financial data plausible, but lacks sender/context.
Fax via Voicemail Office Easy 13 2 1 Low alignment: Fax delivery uncommon in modern workplace.
Email Not Delivered — Office Easy 13 2 3 High alignment: Bounce messages routine and widely expected.
Email Not Delivered — Google Easy 16 3 3 High alignment: Gmail bounce notices with technical detail believable.
From Marketing: Collab Request Easy 14 2 3 High alignment: Internal collaboration messages from marketing common.
Accounting Team Amex Charges — Office Easy 16 3 3 High alignment: Finance inquiries during month-end familiar.
HSA Use Balance! Easy 17 3 2 Moderate alignment: HSA reminders expected, but lacks personalization.
Tax Workshop Easy 16 3 2 Moderate alignment: HR tax webinars plausible but weakened by generic formatting.
Device Non-compliant - Microsoft Medium 15 3 3 High alignment: Matches expected MDM behavior and terminology.
Mobile Device Restricted — Outlook Medium 16 3 3 High alignment: Device quarantine messages match common security flows.
Device Non-compliant - Google Medium 16 3 3 High alignment: Technical restriction notices align with enterprise environments.
Attachment - Word Hard 14 2 2 Moderate alignment: Document sharing plausible, but placeholder fields reduce realism.
Vulnerabilities — Office Hard 18 3 1 Low alignment: Scenario vague and not role-specific; unlikely context.
Loom - Recording Shared Hard 14 2 2 Moderate alignment: Loom recordings common, but template fields weaken trust.
Mentor Program Hard 15 3 3 High alignment: Internal mentorship invitations believable and well-framed.
HR - Performance Review — Outlook Hard 16 3 3 High alignment: Outlook calendar invites for reviews routine.
Changes to Healthcare Form Hard 15 3 3 High alignment: Healthcare form updates common in HR communications.
Vulnerabilities — Google Hard 16 3 1 Low alignment: Scenario lacks specificity and role-relevant context.
Appendix C Lures
See Table 6 for our analysis of the phishing lures used in this study.
Following the NIST Phish Scale framework, we created a 3×3 classification matrix combining cue visibility (few-low, some-medium, many-high) with premise alignment (weak-low, medium, strong-high). Templates were assigned composite difficulty scores by summing both dimensions, with scores ranging from 2 (easy: high cues, low alignment) to 6 (hard: low cues, high alignment). Based on lure and premise scores, we mapped the top group to “Easy”, the middle group to “Medium”, and the bottom group to “Hard”.
The deployed templates were constrained by organizational considerations. ACME Corp.’s management requested the exclusion of highly distressing content (e.g., emergency family situations, personnel security scenarios) to minimize employee psychological impact. This constraint reflects the practical limitations of conducting research within operational environments, where employee welfare considerations influence experimental designs [25].